# Broad Instantaneous Bandwidth Microwave Spectrum Analyzer with a Microfabricated Atomic Vapor Cell

### v1

@@ IV.0.2 Instantaneous bandwidth
A key performance metric of our atomic MW spectrum analyzer is the instantaneous bandwidth $\Delta\omega_{\mathrm{BW}}$, which is primarily determined by the magnetic field gradient and cell length, see Eq. (3). We measure $\Delta\omega_{\mathrm{BW}}$ for our three MEMS atomic vapor cells of different length and the results are presented in Table 1. The exact positions of the three cells with respect to the permanent magnets are not identical, resulting in slightly different magnetic field gradients. The longest MEMS cell with $L=10$ mm reaches a maximal instantaneous bandwidth of $\Delta\omega_{\mathrm{BW}}/(2\pi)\approx 1.0$ GHz. Note that the frequency resolution values for the $L=6$ mm and $L=10$ mm cells may not be optimal since the MW pulse duration was not adjusted. To further increase the instantaneous bandwidth, both the magnetic field gradient and the cell length can be increased.
@@
Figure 4: (a) Images of $\Delta\textrm{OD}_{\textrm{MW}}$ for single-tone MW signals of different power $P_{\mathrm{MW}}$. Here, $\tau_{\mathrm{MW}}=2~{}\mu\mathrm{s}$ and other conditions are as in Fig. 2(a). (b) Peak values of the Gaussian fit to the integrated spectrum $\Sigma_{y}(\Delta\mathrm{OD}_{\mathrm{MW}})$ with $y\in\left[0,1.2\,\textrm{mm}\right]$, as a function of $P_{\mathrm{MW}}$. The linear fit is performed on the leftmost 9 points, the magenta dashed line is the noise level without MW, and the blue solid line is the photon shot noise level.
@@ V Theoretical simulations
We simulate the atomic spectrum analyzer images in order to understand the physical processes in greater detail and to quantitatively compare with the data. Our numerical simulation models the dynamics of a ${}^{87}$b atomic vapor in a two-dimensional, thin cell, taking into account the Hamiltonian of all 8 levels of the $5^{2}\mathrm{S}_{1/2}$ ground state and all 16 levels of the $5^{2}\mathrm{P}_{3/2}$ excited state [50]. We consecutively model the steps of our measurement sequence, consisting of optical pumping, microwave interaction, and laser absorption imaging, taking into account atomic diffusion and collisional relaxation.
Optical pumping. A thorough description of optical pumping in atomic vapors with diffusing atoms is given in [51]. There, the equations of motion are solved in Liouville space, while here we rely on a rate equation approach [52]. By assuming that the optical coherences of the density matrix of the ${}^{87}$b atoms quickly reach a steady state, and the decay rates from the excited states are much higher than the excitation rates, we can describe the optical pumping process with partial differential equations for the populations $p_{n}\equiv p_{n}(t,z,y)$, where $n=1,...,8$:
@@
VI Discussion
We discuss possible improvements to our atomic spectrum analyzer based on theoretical simulations with the model described in the previous section. To increase the instantaneous bandwidth while maintaining a practical cell length, a larger static field gradient is required. Considering a cell with $L=2$ cm and $t_{1}=675~{}\mu$m thickness in a realistic magnetic field gradient of $G=50~{}\mathrm{T}/\mathrm{m}$, the instantaneous frequency span exceeds $25$ GHz. Using our theoretical simulation, we calculate the dependence of the frequency resolution $\gamma/(2\pi)$ and MW power sensitivity (normalized to $1$ s integration time) on MW pulse length $\tau_{\mathrm{MW}}$ and buffer gas pressure $P_{\mathrm{N_{2}}}$, see Fig. 7. For each set of parameters, we employ an optimized algorithm to enhance SNR for weak MW signals by selecting an effective image region. MW power sensitivity is defined as the equivalent photon-shot noise level over the selected pixels at optimal probe power.
Fig. 7 shows that higher buffer gas pressures improve frequency resolution by decreasing diffusion distance, but at the same time decrease sensitivity, because larger collisional broadening decreases the initial optical depth $\mathrm{OD}$ as well as the value after optical pumping $\Delta\mathrm{OD}_{\mathrm{pump}}$. Longer MW pulses improve sensitivity but also broaden frequency resolution (compare Fig. 3(b)). For $\tau_{\mathrm{MW}}=0.8~{}\mu$s and $150$ mbar $\mathrm{N}_{2}$ buffer gas, a frequency resolution of $50$ MHz is possible, meaning that $\approx 500$ frequency channels can be simultaneously detected over the $25$ GHz span, with single-channel power sensitivity of $-55$ dBm.
@@ VI Discussion
Real-time operation is a crucial capability for spectrum analyzers, promising gap-less coverage of all incident broadband MW events in the time domain. Our theoretical model suggests that an efficiently optically pumped steady state can be established within microseconds with higher pump power. Therefore, operating multiple atomic spectrum analyzers simultaneously allows for the design of experimental sequences to consecutively allocate their MW detection windows in the time domain, achieving a high probability of interception for rapidly changing events.

### v2

@@ IV.0.2 Instantaneous bandwidth
A key performance metric of our atomic MW spectrum analyzer is the instantaneous bandwidth $\Delta\omega_{\mathrm{BW}}$, which is primarily determined by the magnetic field gradient and cell length, see Eq. (3). We measure $\Delta\omega_{\mathrm{BW}}$ by applying two microwave tones corresponding to the left and right edge of the cell, respectively. The results for our three MEMS atomic vapor cells of different length are presented in Table 1.
Figure 4: (a) Measurement of $\Delta\omega_{\mathrm{BW}}$ with the $L=10$ mm cell. The gray dashed lines indicate the cell edges. Here, $\tau_{\mathrm{MW}}=2~{}\mu\mathrm{s}$ and the two microwave tones corresponding to the left and right cell edge are $\omega_{\mathrm{MW}}/(2\pi)=12.91~{}\mathrm{GHz}$ and $\omega_{\mathrm{MW}}/(2\pi)=13.88~{}\mathrm{GHz}$, respectively. The two tones are applied consecutively and the images are stitched together at $z=0$. (b) MW spectrum obtained by integrating the image in (a) over $y\in\left[0,1.2~{}\textrm{mm}\right]$, with Gaussian fits to the two MW signals close to the cell edges.
@@
Figure 5: (a) Images of $\Delta\textrm{OD}_{\textrm{MW}}$ for single-tone MW signals of different power $P_{\mathrm{MW}}$. Here, $\tau_{\mathrm{MW}}=2~{}\mu\mathrm{s}$ and other conditions are as in Fig. 2(a). (b) Peak values of the Gaussian fit to the integrated spectrum $\Sigma_{y}(\Delta\mathrm{OD}_{\mathrm{MW}})$ with $y\in\left[0,1.2\,\textrm{mm}\right]$, as a function of $P_{\mathrm{MW}}$ (error bars: 1 SE confidence interval). The linear fit is performed on the leftmost 9 points, the magenta dashed line is the noise level without MW, and the blue solid line is the photon shot noise level.
The exact positions of the three cells with respect to the permanent magnets are not identical, resulting in slightly different magnetic field gradients. The longest MEMS cell with $L=10$ mm reaches a maximal instantaneous bandwidth of $\Delta\omega_{\mathrm{BW}}/(2\pi)\approx 1.0$ GHz, see Fig. 4. Note that the frequency resolution values for the $L=6$ mm and $L=10$ mm cells may not be optimal since the MW pulse duration was not adjusted. To further increase the instantaneous bandwidth, both the magnetic field gradient and the cell length can be increased.
@@ V Theoretical simulations
We simulate the atomic spectrum analyzer images in order to understand the physical processes in greater detail and to quantitatively compare with the data. Our numerical simulation models the dynamics of a 87Rb atomic vapor in a two-dimensional, thin cell, taking into account the Hamiltonian of all 8 levels of the $5^{2}\mathrm{S}_{1/2}$ ground state and all 16 levels of the $5^{2}\mathrm{P}_{3/2}$ excited state [51]. We consecutively model the steps of our measurement sequence, consisting of optical pumping, microwave interaction, and laser absorption imaging, taking into account atomic diffusion and collisional relaxation.
Optical pumping. A thorough description of optical pumping in atomic vapors with diffusing atoms is given in [52]. There, the equations of motion are solved in Liouville space, while here we rely on a rate equation approach [53]. By assuming that the optical coherences of the density matrix of the 87Rb atoms quickly reach a steady state, and the decay rates from the excited states are much higher than the excitation rates, we can describe the optical pumping process with partial differential equations for the populations $p_{n}\equiv p_{n}(t,z,y)$, where $n=1,...,8$:
@@
VI future improvements
We discuss possible improvements to our atomic spectrum analyzer based on theoretical simulations with the model described in the previous section. To increase the instantaneous bandwidth while maintaining a practical cell length, a larger static field gradient is required. Considering a cell with $L=2$ cm and $t_{1}=675~{}\mu$m thickness in a realistic magnetic field gradient of $G=50~{}\mathrm{T}/\mathrm{m}$ [57], the instantaneous frequency span exceeds $25$ GHz. Using this gradient and cell dimension in our theoretical simulation, we calculate the dependence of the frequency resolution $\gamma/(2\pi)$ and MW power sensitivity (normalized to $1$ s integration time) on MW pulse length $\tau_{\mathrm{MW}}$ and buffer gas pressure $P_{\mathrm{N_{2}}}$, see Fig. 8. For each set of parameters, we employ an optimized algorithm to enhance SNR for weak MW signals by selecting an effective image region. MW power sensitivity is defined as the equivalent photon-shot noise level over the selected pixels at optimal probe power.
Fig. 8 shows that higher buffer gas pressures improve frequency resolution by decreasing diffusion distance, but at the same time decrease sensitivity, because larger collisional broadening decreases the initial optical depth $\mathrm{OD}$ as well as the value after optical pumping $\Delta\mathrm{OD}_{\mathrm{pump}}$. Longer MW pulses improve sensitivity but also reduce frequency resolution (compare Fig. 3(b)). For $\tau_{\mathrm{MW}}=0.8~{}\mu$s and $150$ mbar $\mathrm{N}_{2}$ buffer gas, a frequency resolution of $50$ MHz is possible, meaning that $\approx 500$ frequency channels can be simultaneously detected over the $25$ GHz span, with single-channel power sensitivity of $-55$ dBm.
@@
Real-time operation is a crucial capability for spectrum analyzers, promising gap-less coverage of all incident broadband MW events in the time domain. Our theoretical model suggests that an efficiently optically pumped steady state can be established within microseconds with higher pump power. Therefore, operating multiple atomic spectrum analyzers simultaneously allows for the design of experimental sequences to consecutively allocate their MW detection windows in the time domain, achieving a high probability of interception for rapidly changing events. Exploiting the fact that the atoms do not significantly attenuate the analyzed MW signal, multiple atomic spectrum analyzers can be connected in series for real-time operation or even broader spectral coverage. This would turn the anticipated $25$ GHz instantaneous bandwidth into a corresponding real-time analysis bandwidth.
Although our atomic spectrum analyzer is still in a proof-of-principle development stage, it is interesting to compare the present and anticipated performance with state-of-the-art electronic spectrum analyzers. Commercial real-time spectrum analyzers currently offer up to 2 GHz real-time bandwidth. These analyzers rely on time-overlapped FFT configurations, requiring powerful digital processing engines and large memory capacities. Users can choose from several options for frequency resolution (i.e. resolution bandwidth) within a given real-time span. This selection impacts the minimum measurement duration (time resolution) and noise level. Finer frequency resolution requires longer measurement durations but achieves lower noise levels. Typical values for frequency resolution range from sub-MHz to about 100 MHz, with minimum measurement durations from several hundreds of microseconds to several microseconds. The noise level for each resolvable frequency channel correspondingly ranges from -110 dBm to -60 dBm. In our current demonstration, the atomic spectrum analyzer already reaches $1$ GHz instantaneous bandwidth, while the improvements outlined above will make it possible to reach $25$ GHz real-time bandwidth at a resolution of $50$ MHz. The minimum measurement duration is a few microseconds. Our experimental demonstration has a sensitivity of $-23$ dBm in $1$ s integration time, while for the improved design we expect a sensitivity of $-55$ dBm in $1$ s integration time, which could be enhanced further by improving the optical probing of the atoms. The main promise of our approach is thus in increasing the real-time bandwidth beyond the capabilities of electronic spectrum analyzers, while only minimally perturbing the signal so that it can be further used while being analyzed.